OpenStudy (anonymous):

Did I do this wrong? Given a cylinder with a surface area of 60 units squared and a height of 7 units. Find the volume of a sphere that has the same size radius as the given cylinder. Show all work. 2πr^2+2πrh = 60 pi 2pi(r^2 + rh)=60pi r^2+r•h=30 h =7 r^2+7•r=30 r(r+7) =30 if r = 3 then 3(3+7)=30 √ r=3

OpenStudy (anonymous):

I think so. I got the same answer: r=3 just put in formula for the volume of a sphere.

OpenStudy (anonymous):

then what do i do next? i dont think im finished with the equation.

OpenStudy (anonymous):

You have to find the volume of the sphere that has that radius.

OpenStudy (anonymous):

though, I solve quadratic equation for r by using quadratic formula. answer= the same

OpenStudy (anonymous):

V=4/3 pi*r put r=3 in a formula

OpenStudy (anonymous):

So just plug it into the equation for the volume of a sphere.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

is it 4/3πr^3 or just 4/3πr?

OpenStudy (anonymous):

I think u need find radius

OpenStudy (anonymous):

you already have radios: r=3 from the previous calculations

OpenStudy (anonymous):

It's \[V = \frac{4}{3}\pi r^3\]